A wire 48 centimeters long is cut into two pieces. Each piece is bent to form a square. Where should the wire be cut so that the sum of the areas of the squares is equal to 1280 square centimeters?
Answer by Paul(988)

Let one peice be x
Let other pice be y

x+y=48
y=48-x (subsitution)

x^2+y^2=1280 (subsitute)
x^2+(48-x)^2=1280
x^2+2304-96x+x^2=1280
2x^2-96x+1024=0,x
Factor
x=16 and x=32
y=48-32
y=16

Hence, the 2 pices are 16cm and 32cm.
Paul.